# Hamba: Single-view 3D Hand Reconstruction with Graph-guided Bi-Scanning Mamba

### v1

@@ Abstract
3D Hand reconstruction from a single RGB image is challenging due to the articulated motion, self-occlusion, and interaction with objects. Existing SOTA methods employ attention-based transformers to learn the 3D hand pose and shape, but they fail to achieve robust and accurate performance due to insufficient modeling of joint spatial relations. To address this problem, we propose a novel graph-guided Mamba framework, named Hamba, which bridges graph learning and state space modeling. Our core idea is to reformulate Mamba’s scanning into graph-guided bidirectional scanning for 3D reconstruction using a few effective tokens. This enables us to learn the joint relations and spatial sequences for enhancing the reconstruction performance. Specifically, we design a novel Graph-guided State Space (GSS) block that learns the graph-structured relations and spatial sequences of joints and uses 88.5% fewer tokens than attention-based methods. Additionally, we integrate the state space features and the global features using a fusion module. By utilizing the GSS block and the fusion module, Hamba effectively leverages the graph-guided state space modeling features and jointly considers global and local features to improve performance. Extensive experiments on several benchmarks and in-the-wild tests demonstrate that Hamba significantly outperforms existing SOTAs, achieving the PA-MPVPE of 5.3mm and F@15mm of 0.992 on FreiHAND. Hamba is currently Rank 1 in two challenging competition leaderboards [^1] HO3Dv2 and HO3Dv3 Competition Leaderboards: https://codalab.lisn.upsaclay.fr/competitions/4318#results, https://codalab.lisn.upsaclay.fr/competitions/4393#results on 3D hand reconstruction. The code will be available upon acceptance.
@@ 1 Introduction
3D Hand reconstruction has numerous applications across various fields including robotics, animation, human-computer interaction, and AR/VR [11, 32, 65]. However, reconstructing 3D hands from a single RGB image without body context or camera parameters remains a difficult challenge in computer vision. Recent works primarily explored transformers [80] for this task. Transformer-based methods [14, 17, 47, 48, 64, 88] have achieved SOTA performance by utilizing attention strategies. METRO [47] introduced a multi-layer transformer, primarily using self-attention and learning vertex-vertex and vertex-joint relations. MeshGraphormer [48] integrated graph convolutions with a transformer to further enhance the reconstruction performance. Recently, HaMeR [64] designed a ViT-based model [17], using ViTPose [88] weights and large datasets to achieve better performance.
However, the above models fail to reconstruct a robust mesh in challenging in-the-wild scenarios that have occlusions, truncation, and hand-hand or hand-object interactions (See Figure 5 for visual comparison). This is partially due to a lack of accurately modeling spatial relations between hand joints. Secondly, transformer-based methods [17, 47, 48, 64, 88] require a large number of tokens for reconstruction, and applying attention to all image tokens does not fully utilize the joint spatial sequences, which often results in an inaccurate 3D hand mesh in real-world scenarios.
To address these challenges, we propose Hamba, a novel Mamba-based [24] model that employs graph learning [41, 96] and state space modeling [24] for robust 3D hand mesh reconstruction. Mamba is a new state space modeling method with global receptive field capability. Most Mamba-based models [3, 24, 31, 45, 81, 82, 95] are designed for long-range data, and very few studies [44, 71] have adapted Mamba for 3D vision tasks. In this work, we explore Mamba’s potential for the 3D hand reconstruction task. We found that directly applying Mamba for 3D hand reconstruction results in inaccurate meshes due to its unidirectional scanning and the lack of specific design for 3D hand reconstruction. To tackle this challenge, we propose a Graph-guided Bidirectional Scan (GBS) to effectively capture the semantic and spatial relation between joints, as shown in Figure 2(c). Besides, the transformer’s attention requires calculating correlation among all tokens and introduces unnecessary background correlations, while our proposed GBS uses 88.5% fewer tokens (see Section 3.2 for more details). Secondly, though Mamba-based models [3, 24, 31, 45, 81, 95] excel in modeling long-range sequences, they are not proficient at capturing the local-relation information (in our case, the ‘semantics’ between hand joints). Since graph learning has the capability to effectively capture node relations, we integrate graph convolutions into state space modeling, significantly enhancing the representation by considering the intricate hand joint relations.
@@
In particular, to effectively leverage state space modeling (SSM) and graph learning capabilities for 3D hand reconstruction, we first carefully design a Token Sampler (TS) under guidance with hand joints predicted by Joint Regressor (JR), then feed sampled token into the Graph-guided State Space (GSS) block under the Graph-guided Bidirectional Scan (GBS). Lastly, we introduce a fusion module to integrate the state space tokens and global features to further improve performance. As shown in Figure 1, Hamba achieves significant visual performance in challenging scenarios. We summarize our contributions as follows:
@@ 3.1 Preliminaries
Figure 3: Overview of Hamba’s model architecture. Given a hand image $I$, we extract the tokens via the backbone model and downsample layers. We design a graph-guided SSM as a decoder to regress hand parameters. Here, the hand joints ($J_{\text{2D}}$) are regressed by the Joints Regressor (JR) which is then input to the Token Sampler (TS) to sample tokens ($T_{\text{TS}}$). The joint spatial sequence tokens ($T_{\text{GSS}}$) are learned by the Graph-guided State Space (GSS) blocks. Note that GCN takes $T_{\text{TS}}$ as input while its output is concatenated with the mean of downsampled tokens. GSS leverages graph learning and state space modeling to capture the joint spatial relations to achieve robust 3D reconstruction.
@@ 3.2 Hamba
Graph-guided Bidirectional Scan (GBS). To achieve robust reconstruction and leverage effective tokens, we reformulate Mamba’s unidirectional scanning as a graph-guided bidirectional scan, thus adapting it for 3D reconstruction tasks. GBS is designed to follow a specific graph pattern, considering the spatial and topological connection of the hand joints with image features. A naive approach would be scanning all tokenized image patches (Figure 2(b)). However, this involves redundant tokens, making it challenging to learn joint spatial relations effectively. To address this, we propose two novel ideas. First, instead of scanning all tokens unidirectionally, we perform hand joint-level bidirectional scanning of sampled tokens $T_{\text{TS}}$. This effectively reduces the number of tokens to be scanned from 192 to 22 ($\approx$ 88.5% reduction). We adapt VMamba [51]’s SS2D block for bidirectional scanning to be suitable for our joint spatial sequence. Second, to capture the local and global joint relations, we introduce a Semantic GCN block [97]. Mamba learns long-range dependencies, but it is less effective at capturing fine-grained local information in complex structures like the 3D mesh. The GCN learns input-independent weight matrix for hand joint edges representing hand priors implied in graph structures, i.e., how one hand joint influences another. Introducing graph learning makes it possible to explicitly encode the graph structure within our GSS module. Let $\mathcal{G}={\mathbf{\{V,E\}}}$ be the graph, $\mathbf{V}$ is the set of $J$ nodes and $\mathbf{E}$ are the edges. $T_{\text{GCN}_{l}}$ represents the output of the $l$-th GCN block, while the complete output of the GSS block is $T_{\text{GSS}_{l}}$. For a graph-based propagation, we multiply the token with a learnable parameter matrix $\mathbf{W}\in\mathbb{R}^{C\times C}$. Thus, the GCN operation is formulated as,
@@
where $\oplus$ denotes concatenation. The GSS block not only leverages features from state space modeling and graph learning but also considers global features. This design enables our Hamba to learn effective features to enhance performance by incorporating state space modeling and graph learning with few tokens, shown in our ablation study Section 4.2.
State Space Modeling for Joint Spatial Sequence. Different from the video-based Mamba models [7, 20, 42] learning the temporal feature with the frame sequence, our Hamba focuses on the joint spatial sequence per frame and reveals that modeling joint relations with Mamba [24] can significantly improve the 3D reconstruction performance. In particular, as shown in Eq. 1, $x(t)$ represents $t$-th token of the joint spatial sequence, which is first sampled by the TS using the JR and then encoded with the GCN. Note that $t$ denotes the index of the hand joint iteration. Lastly, $y(t)$ is the updated token of the $t$-th of the joint spatial sequence after passing through GSS Blocks. The proposed GSS block effectively enhances 3D reconstruction performance by learning the joint spatial sequence relations with graph learning and state space modeling.
Loss Functions. Following [64], we train Hamba using a combined loss which includes 2D joint loss $\mathcal{L}_{\text{2D}}$, 3D joint loss $\mathcal{L}_{\text{3D}}$, pose loss $\mathcal{L}_{\theta}$, shape loss $\mathcal{L}_{\beta}$, and an adversarial loss $\mathcal{L}_{\text{adv}}$. Here, $\mathcal{L}_{\text{2D}}$ and $\mathcal{L}_{\text{3D}}$ are calculated using the L1 Norm, while $\mathcal{L}_{\theta}$ and $\mathcal{L}_{\beta}$ are computed using the L2 Norm. The training loss $\mathcal{L}_{\text{total}}$ is defined as,
@@
where $\lambda_{\text{2D}},~{}\lambda_{\text{3D}},~{}\lambda_{\theta},~{}\lambda_{\beta}$, and $\lambda_{\text{adv}}$ denote the weights of $\mathcal{L}_{\text{2D}},~{}\mathcal{L}_{\text{3D}},~{}\mathcal{L}_{\theta},~{}\mathcal{L}_{\beta}$, and $\mathcal{L}_{\text{adv}}$, respectively.
@@ 4.1 Main Results
In-the-wild Generalizability Evaluation. Approximately $95\%$ of datasets used for training previous models [9, 14, 43, 47, 48, 64, 79, 98] were collected in controlled indoor environments, such as studios or multi-camera setups. This includes the FreiHAND [102], HO3Dv2 [27], and HO3Dv3 [28] benchmarks that are popularly used for both training and evaluation. However, training models on datasets collected in controlled environments often leads to decreased performance in real-world scenarios. Thus, solely evaluating performance over indoor-collected datasets might not provide a correct evaluation of the robustness of 3D hand reconstruction. We additionally evaluate Hamba’s in-the-wild performance on the recently proposed HInt [64] benchmark, which has variations in visual conditions, viewpoints, and hand interactions. Since the annotations of the HInt-NewDays [13] and HInt-EpicKitchensVISOR [16] benchmark dataset are 2D keypoints, PCK [90] computed at varying thresholds is used as the evaluation metrics. As shown in Table 4, Hamba outperforms existing models by a large margin and surpasses HaMeR [64], showing the improvement in model robustness for in-the-wild scenarios. None of the models (including Hamba) have been trained on/ previously ever seen the HInt dataset.
Figure 5: Qualitative in-the-wild comparison of the proposed Hamba with SOTAs on HInt-EpicKitchensVISOR [16, 64]. No models (including Hamba) have been trained on HInt.
@@
Qualitative Comparison. Figure 5 presents the qualitative comparison of Hamba’s 3D hand mesh reconstruction with SOTA models on in-the-wild images from HInt-EpicKitchens. This includes models that directly regress vertices (METRO [47], MeshGraphormer [48]), and parametric methods, which regress MANO parameters (FrankMocap [69], HaMeR [64]). These images are particularly challenging since they comprise real-world cooking videos of a person with highly occluded hands, hand-hand, and/or hand-object interactions. For visual comparison, we select images where the hand lies in the corners, causing a truncation scenario thus increasing the complexity further. Hamba consistently outperforms other models and achieves a much better reconstruction. From Figure 5, we can observe that in severe in-the-wild truncation scenarios, Hamba achieves better hand reconstruction, even though the hand is truncated or occluded. We attribute this performance to effectively learning the spatial hand joint sequence with the state space model. The same is verified in the ablation study presented in Sec. 4.2. Figure 7 presents in-the-wild results on various movies, interviews, etc., scenarios. Figure 8 and Figure 9 presents additional visual results on HInt-NewDays and HInt-EpicKitchensVISOR respectively. Hamba can robustly reconstruct 3D hands in various complicated hand gestures like grasping, holding, grabbing, finger-pointing, and flattening from different viewing directions, even in heavily occluded and truncated scenarios.
Table 2: Comparison with SOTAs on HO3Dv2 [27] hand-object interaction benchmark. Backbones are the same as in Table 1.
@@ 4.2 Ablation Studies
Effect of Branch-wise Features. We verify the effectiveness of each branch feature by excluding their respective tokens from the fusion module as shown in Table 5. First, we verify the contribution of the proposed GSS branch. When the GSS tokens are excluded (Row 3), we observe a major drop in model performance. Specifically, F@5mm ($\uparrow$) drops from 0.738 $\rightarrow$ 0.717, and the PA-MPJPE ($\downarrow$) and PA-MPVPE ($\downarrow$) errors increase from 6.6 $\rightarrow$ 6.9 and 6.3 $\rightarrow$ 6.6. Thus, in addition to local and global contexts, incorporating structured state-space representations can be effective for 3D hand reconstruction. Moreover, it is important to note that modeling spatial joint sequence relations provides better tokens than directly using the 2D joint locations, even though the latter has a clear semantic meaning for all the hand joints. We attribute this to cases of occlusions where the 2D joints cannot be precisely predicted. Removing the Token sampler (Row 1) or the 2D joints (Row 2) features also shows a performance drop, but is less significant than removing the GSS branch, since they only provide the local context while GSS tokens provide both local and spatial-relations information. Note that the Global Mean token (Row 4) remains important since it captures the global context, which is discarded in the ablation.
Effect of proposed Components. Since the GSS Block stands as a major contribution, we additionally evaluate the effectiveness of each component in the proposed GSS block. The same is presented in Table 5. The GSS block models the hand-joint topological structure, learning the graph-structured relations and spatial sequences of joints via graph and state space modeling. Adopting graph learning additionally provides the local context. Excluding the GCN, i.e., when simply using a Mamba block, the structure information will be neglected from the input to the SS2D block, which leads to a large drop in performance (Row 7). This indicates that the GCN is an essential component of the GSS block and using SS2D blocks alone does not lead to accurate 3D hand mesh reconstruction. A potential counter-argument may be that the input features and the GCN alone are sufficient for 3D hand reconstruction, without much improvement from the Mamba Blocks. We removed the Mamba blocks and the GSS degenerates into simple GCN, leading to an equal performance drop (Row 9). Specifically, the PA-MPJPE ($\downarrow$) and PA-MPVPE ($\downarrow$) increase from 6.6 $\rightarrow$ 7.3 and 6.3 $\rightarrow$ 7.2 respectively, while the F@5mm ($\uparrow$) and F@15mm ($\uparrow$) drop from 0.738 $\rightarrow$ 0.675 and 0.988 $\rightarrow$ 0.983 respectively. This confirms that both the GCN and the Mamba blocks are equally important in the GSS Block. To verify the effectiveness of the bidirectional scanning, we replaced it with conventional unidirectional scanning to compare, denoted as w/o Bidirectional-scan (Row 6), and the reconstruction error increased. An even larger drop in performance is observed when the proposed Graph-guided bidirectional scan is completely removed from the model (Row 8). When not using the token sampler, we also see a drop in performance (Row 5). Overall, Table 5 verifies the effectiveness of each proposed component. We additionally validate this by a qualitative evaluation (in Figure 6), wherein the visual result gets worse when we remove our proposed modules.
Table 4: In-the-wild Generalizability Evaluation results on the HInt [64] benchmark. Here, PCK is used as the evaluation metric.
Method Venue NewDays VISOR
@0.05$\uparrow$ @0.1$\uparrow$ @0.15$\uparrow$ @0.05$\uparrow$ @0.1$\uparrow$ @0.15$\uparrow$
All Joints METRO [47] CVPR 21 14.7 38.8 57.3 16.8 45.4 65.7
FrankMocap [69] ICCVW 21 16.1 41.4 60.2 16.8 45.6 66.2
MeshGraphormer [48] ICCV 21 16.8 42.0 59.7 19.1 48.5 67.4
HandOccNet (param) [62] CVPR 22 9.1 28.4 47.8 8.1 27.7 49.3
HandOccNet (no param) [62] CVPR 22 13.7 39.1 59.3 12.4 38.7 61.8
HaMeR [64] CVPR 24 48.0 78.0 88.8 43.0 76.9 89.3
Hamba Ours 48.7 79.2 90.0 47.2 80.2 91.2
Visible Joints METRO [47] CVPR 21 19.2 47.6 66.0 19.7 51.9 72.0
FrankMocap [69] ICCVW 21 20.1 49.2 67.6 20.4 52.3 71.6
Mesh Graphormer [48] ICCV 21 22.3 51.6 68.8 23.6 56.4 74.7
HandOccNet (param) [62] CVPR 22 10.2 31.4 51.2 8.5 27.9 49.8
HandOccNet (no param) [62] CVPR 22 15.7 43.4 64.0 13.1 39.9 63.2
HaMeR [64] CVPR 24 60.8 87.9 94.4 56.6 88.0 94.7
Hamba Ours 61.2 88.4 94.9 61.4 89.6 95.6
Occluded Joints METRO [47] CVPR 21 7.0 23.6 42.4 10.2 32.4 53.9
FrankMocap [69] ICCVW 21 9.2 28.0 46.9 11.0 33.0 55.0
MeshGraphormer [48] ICCV 21 7.9 25.7 44.3 10.9 33.3 54.1
HandOccNet (param) [62] CVPR 22 7.2 23.5 42.4 7.4 26.1 46.7
HandOccNet (no param) [62] CVPR 22 9.8 31.2 50.8 9.9 33.7 55.4
HaMeR [64] CVPR 24 27.2 60.8 78.9 25.9 60.8 80.7
Hamba Ours 28.2 62.8 81.1 29.9 66.6 84.3
@@
Table 5: Ablation study on the FreiHAND dataset [102] to verify the effect of proposed components. All variants are trained for the same number of steps. Without is abbreviated as ‘w/o’.
Ablation PA-MPJPE $\downarrow$ PA-MPVPE $\downarrow$ F@5mm $\uparrow$ F@15mm $\uparrow$

### v2

@@ Abstract
3D Hand reconstruction from a single RGB image is challenging due to the articulated motion, self-occlusion, and interaction with objects. Existing SOTA methods employ attention-based transformers to learn the 3D hand pose and shape, yet they do not fully achieve robust and accurate performance, primarily due to inefficiently modeling spatial relations between joints. To address this problem, we propose a novel graph-guided Mamba framework, named Hamba, which bridges graph learning and state space modeling. Our core idea is to reformulate Mamba’s scanning into graph-guided bidirectional scanning for 3D reconstruction using a few effective tokens. This enables us to efficiently learn the spatial relationships between joints for improving reconstruction performance. Specifically, we design a Graph-guided State Space (GSS) block that learns the graph-structured relations and spatial sequences of joints and uses 88.5% fewer tokens than attention-based methods. Additionally, we integrate the state space features and the global features using a fusion module. By utilizing the GSS block and the fusion module, Hamba effectively leverages the graph-guided state space features and jointly considers global and local features to improve performance. Experiments on several benchmarks and in-the-wild tests demonstrate that Hamba significantly outperforms existing SOTAs, achieving the PA-MPVPE of 5.3mm and F@15mm of 0.992 on FreiHAND. At the time of this paper’s acceptance, Hamba holds the top position, Rank 1, in two competition leaderboards [^1] HO3Dv2 and HO3Dv3 Leaderboards: https://codalab.lisn.upsaclay.fr/competitions/4318#results, https://codalab.lisn.upsaclay.fr/competitions/4393#results on 3D hand reconstruction.
11footnotetext: Equal contribution. Corresponding author.
@@ 1 Introduction
3D Hand reconstruction has numerous applications across multiple fields, which include robotics, animation, human-computer interaction, and AR/VR [11, 34, 71, 18, 103]. However, reconstructing 3D hands from a single RGB image without body context or camera parameters remains a difficult challenge in computer vision. Recent works primarily explored transformers [14, 19, 51, 52, 70, 95, 73, 45, 55] for this task and achieved SOTA performance by utilizing attention mechanism. METRO [51] introduced a multi-layer transformer, using self-attention to learn vertex-vertex and vertex-joint relations. MeshGraphormer [52] integrated graph convolutions with a transformer to further enhance the reconstruction performance. Recently, HaMeR [70] designed a ViT-based model [19], using ViTPose [95] weights and large datasets to achieve better performance.
However, the above models fail to reconstruct a robust mesh in challenging in-the-wild scenarios that have occlusions, truncation, and hand-hand or hand-object interactions (See Figure 5 for visual comparison). This is partially due to a lack of accurate modeling of spatial relations among hand joints. Secondly, transformer-based methods [19, 51, 52, 70, 95] require a large number of tokens for reconstruction, and applying attention to all image tokens does not efficiently model the joint spatial sequences (i.e., the spatial relationship between joints), which often results in an inaccurate 3D hand mesh in real-world scenarios.
@@
To address these challenges, we propose Hamba, a novel Mamba-based [26] model that employs graph learning [44, 104] and state space modeling [26] for robust 3D hand mesh reconstruction. Mamba is a new state space modeling method with global receptive field capability. Most Mamba-based models [3, 26, 33, 49, 88, 89, 102] are designed for long-range data, and few studies [48, 78] have adapted Mamba for 3D vision tasks. In this work, we explore Mamba’s potential for the 3D hand reconstruction task. We found that directly applying Mamba for 3D hand reconstruction results in inaccurate meshes due to its unidirectional scanning and the lack of specific design for 3D hand reconstruction. To tackle this challenge, we propose a Graph-guided Bidirectional Scan (GBS) to effectively capture the semantic and spatial relation between joints, as shown in Figure 2(c). Besides, the transformer’s attention requires calculating correlation among all tokens and introduces unnecessary background correlations, while our proposed GBS uses 88.5% fewer tokens (see Section 3.2 for more details). Secondly, though Mamba-based models [3, 26, 33, 49, 88, 102] excel in modeling long-range sequences, they are not proficient at capturing the local-relation information (in our case, the ‘semantics’ between hand joints). Since graph learning has the capability to effectively capture node relations, we integrate graph convolutions into state space modeling, significantly enhancing the representation by considering the intricate hand joint relations.
In particular, to effectively leverage state space modeling (SSM) and graph learning capabilities for 3D hand reconstruction, we first carefully design a Token Sampler (TS) under guidance with hand joints predicted by Joint Regressor (JR), then feed sampled token into the Graph-guided State Space block (GSS) under the Graph-guided Bidirectional Scan (GBS). Lastly, we introduce a fusion module to integrate the state space tokens and global features to further improve performance. As shown in Figure 1, Hamba achieves significant visual performance in challenging scenarios. We summarize our contributions as follows:
@@ 3.1 Preliminaries
Figure 3: Overview of Hamba’s architecture. Given a hand image $I$, tokens are extracted via a trainable backbone model and downsampled. We design a graph-guided SSM as a decoder to regress hand parameters. The hand joints ($J_{\text{2D}}$) are regressed by Joints Regressor (JR) and fed into the Token Sampler (TS) to sample tokens ($T_{\text{TS}}$). The joint spatial sequence tokens ($T_{\text{GSS}}$) are learned by the Graph-guided State Space (GSS) blocks. Inside each GSS block, the GCN network takes $T_{\text{TS}}$ as input and its output is concatenated with the mean down-sampled tokens. GSS leverages graph learning and state space modeling to capture the joint spatial relations to achieve robust 3D reconstruction.
@@ 3.2 Hamba
Graph-guided Bidirectional Scan (GBS). To achieve robust reconstruction and leverage effective tokens, we reformulate Mamba’s unidirectional scanning as a graph-guided bidirectional scan, thus adapting it for 3D reconstruction tasks. GBS is designed to follow a specific graph pattern, considering the spatial and topological connection of the hand joints with image features. A naive approach would be scanning all tokenized image patches (Figure 2(b)). However, this involves redundant tokens, making it challenging to learn joint spatial relations effectively. To address this, we propose two novel ideas. First, instead of scanning all tokens unidirectionally, we perform hand joint-level bidirectional scanning of sampled tokens $T_{\text{TS}}$. This effectively reduces the number of tokens to be scanned from 192 to 22 ($\approx$ 88.5% reduction). We adapt VMamba [57]’s SS2D block for bidirectional scanning to be suitable for our joint spatial sequence. Second, to capture the local and global joint relations, we introduce a Semantic GCN block [105]. Mamba learns long-range dependencies, but it is less effective at capturing fine-grained local information in intricate structures like the 3D mesh. The GCN learns input-independent weight matrix to model the edges between hand joints, reflecting how one joint influences another based on prior embedded in graph structures. Introducing graph learning makes it possible to explicitly encode the graph structure within our GSS module. Let $\mathcal{G}={\mathbf{\{V,E\}}}$ be the graph, $\mathbf{V}$ is the set of $J$ nodes and $\mathbf{E}$ are the edges. $T_{\text{GCN}_{l}}$ represents the output of the $l$-th GCN block, while the complete output of the GSS block is $T_{\text{GSS}_{l}}$. For a graph-based propagation, we multiply the token with a learnable parameter matrix $\mathbf{W}\in\mathbb{R}^{C\times C}$. Thus, the GCN operation is formulated as,
@@
where $\oplus$ denotes concatenation. The GSS block not only leverages features from state space modeling and graph learning but also considers global features. This design enables Hamba to learn effective features to enhance performance by incorporating state space modeling and graph learning with few tokens, shown in our ablation study Section 4.2.
State Space Modeling for Joint Spatial Sequence. Different from the video-based Mamba models [7, 22, 46], which learns the temporal feature with the frame sequence, Hamba focuses on the joint spatial sequence per frame and reveals that modeling joint relations with Mamba [26] can significantly improve the 3D reconstruction performance. In particular, as shown in Eq. 1, $x(t)$ represents $t$-th token of the joint spatial sequence, which is first sampled by the TS using the JR and then encoded with the GCN. Note that $t$ denotes the index of the hand joint iteration. Lastly, $y(t)$ is the updated token of the $t$-th of the joint spatial sequence after passing through GSS Blocks. The proposed GSS block effectively enhances 3D reconstruction performance by learning the joint spatial sequence relations with graph learning and state space modeling.
Loss Functions. Following [70], we train Hamba using a combined loss which includes 2D joint loss $\mathcal{L}_{\text{2D}}$, 3D joint loss $\mathcal{L}_{\text{3D}}$, pose loss $\mathcal{L}_{\theta}$, shape loss $\mathcal{L}_{\beta}$, and an adversarial loss $\mathcal{L}_{\text{adv}}$. $\mathcal{L}_{\text{2D}}$ and $\mathcal{L}_{\text{3D}}$ are calculated using the L1 Norm, while $\mathcal{L}_{\theta}$ and $\mathcal{L}_{\beta}$ use the L2 Norm. The training loss $\mathcal{L}_{\text{total}}$ is defined as Equation 9, where $\lambda_{\text{2D}},~{}\lambda_{\text{3D}},~{}\lambda_{\theta},~{}\lambda_{\beta}$, and $\lambda_{\text{adv}}$ denote each term’s weight respectively.
@@ 4.1 Main Results
In-the-wild Generalizability Evaluation. Approximately $95\%$ of datasets used for training previous models [9, 14, 47, 51, 52, 70, 86, 107] were collected in controlled indoor environments, such as studios or multi-camera setups. This includes the FreiHAND [111], HO3Dv2 [29], and HO3Dv3 [30] benchmarks that are popularly used for both training and evaluation. However, training models on datasets collected in controlled environments often leads to decreased performance in real-world scenarios. Thus, solely evaluating performance over indoor-collected datasets might not provide a correct evaluation of the robustness of 3D hand reconstruction. We additionally evaluate Hamba’s in-the-wild performance on the recently proposed HInt [70] benchmark, which has variations in visual conditions, viewpoints, and hand interactions. Since HInt-NewDays [13] and HInt-EpicKitchensVISOR [16, 17] annotations are 2D keypoints, PCK [97] computed at varying thresholds is used as the evaluation metrics. As shown in Table 5, Hamba outperforms existing models by a large margin and surpasses HaMeR [70], showing improvement in model robustness for in-the-wild scenarios. None of the models (including Hamba) have been trained on/ previously ever seen HInt dataset.
Qualitative Comparison. Figure 5 presents the qualitative comparison of Hamba’s 3D hand mesh reconstruction with SOTA models on in-the-wild images from HInt-EpicKitchens. This includes models that directly regress vertices (METRO [51], MeshGraphormer [52]), and parametric methods, which regress MANO parameters (FrankMocap [76], HaMeR [70]). These images are particularly challenging since they comprise real-world cooking videos of a person with highly occluded hands, hand-hand, and/or hand-object interactions. For visual comparison, we select images where the hand lies in the corners, causing a truncation scenario thus increasing the complexity further. Hamba consistently outperforms other models and achieves a much better reconstruction. From Figure 5, we can observe that in severe in-the-wild truncation scenarios, Hamba achieves better hand reconstruction, even though the hand is truncated or occluded. We attribute this performance to effectively learning the spatial hand joint sequence with the state space model. The same is verified in the ablation study presented in Sec. 4.2. Figure S5 presents in-the-wild results on various movies, interviews, etc., scenarios. Figure S6 and Figure S7 presents additional visual results on HInt-NewDays and HInt-EpicKitchensVISOR respectively. Hamba can robustly reconstruct 3D hands in various complicated hand gestures like grasping, holding, grabbing, finger-pointing, and flattening from different viewing directions, even in heavily occluded and truncated scenarios.
Figure 5: Qualitative in-the-wild comparison of the proposed Hamba with SOTAs on HInt-EpicKitchensVISOR [16, 70]. None of the models (including Hamba) have been trained on HInt.
@@
Table 2: Comparison with SOTAs on HO3Dv2 [29] hand-object interaction benchmark.
@@ 4.2 Ablation Studies
Effect of Branch-wise Features. We verify the effectiveness of each branch feature by excluding their respective tokens from the fusion module as shown in Table 4. First, we verify the contribution of the proposed GSS branch. When the GSS tokens are excluded (Row 3), we observe a major drop in model performance. Specifically, F@5mm ($\uparrow$) drops from 0.738 $\rightarrow$ 0.717, and the PA-MPJPE ($\downarrow$) and PA-MPVPE ($\downarrow$) errors increase from 6.6 $\rightarrow$ 6.9 and 6.3 $\rightarrow$ 6.6. Thus, in addition to local and global contexts, incorporating structured state-space representations can be effective for 3D hand reconstruction. Moreover, it is important to note that modeling spatial joint sequence relations provides better tokens than directly using the 2D joint locations, even though the latter has a clear semantic meaning for all the hand joints. We attribute this to cases of occlusions where the 2D joints cannot be precisely predicted.
@@
Table 4: Ablation study on FreiHAND [111] to verify proposed components. All variants are trained for same number of steps. PA-MPJPE, PA-MPVPE and without are abbreviated as PJ, PV, ‘w/o’.
Ablation PJ $\downarrow$ PV $\downarrow$ F@5 $\uparrow$ F@15 $\uparrow$